State Newton’s laws of motion. Explain how the “newton” is defined from these laws.
1
Expert's answer
2020-11-23T10:29:55-0500

Source: https://en.wikipedia.org/wiki/Newton%27s_laws_of_motion

First law

In an inertial frame of reference, an object either remains at rest or continues to move at a constant velocity, unless acted upon by a force.

Second law

In an inertial frame of reference, the vector sum forces F on an object is equal to the mass m of that object multiplied by the acceleration a of the object: "F = m\times a" . (It is assumed here that the mass m is constant – see below)

Third law

When one body exerts a force on a second body, the second body simultaneously exerts a force equal in magnitude and opposite in direction on the first body.

Since "newton" is "\frac {kg\times m} {s ^ 2}" , and according to Newton's second law, "F = m\times a" therefore "[F] = kg\times\frac {m} {s ^ 2} = N"